VIBRATIONAL SPECTROSCOPIC STUDY ON SOME HOFMANN TYPE CLATHRATES: M(2-(1-CYCLOHEXENYL)ETHYLAMINE)_{2} Ni(CN)_{4}.2BENZENE (M =Ni AND Cd)

TEKİN İZG˙, DEPARTMENT OF PHYSICS, ARTS AND SCIENCE FACULTY, İNÖNÜ UNI-VERSITY, MALATYA, 44069, TURKEY; CEMAL PARLAK, DEPARTMENT OF PHYSICS, ARTS AND SCIENCE FACULTY, DUMLUPINAR UNIVERSITY, KÜTAHYA, 43100, TURKEY; \underline{MUSTAFA SENYEL}, DEPARTMENT OF PHYSICS, SCIENCE FACULTY, ANADOLU UNIVER-SITY, ESKİŞEHİR, 26470, TURKEY.

New Hofmann type benzene clathrates in the form of M(CyHEA)2Ni(CN)4.2Benzene (where CyHEA = 2-(1-Cyclohexenyl)ethylamine and M = Ni or Cd) have been prepared in powder form and FT-IR and Raman spectra have been reported. The results suggest that title compounds are similar in structure to Hofmann type clathrates and their structures consist of polymeric layers of | M-Ni(CN)4 |_{∞} with the CyHEA molecule bounded to the metal atoms (M).